# On-line conformalized neural networks ensembles for probabilistic forecasting of day-ahead electricity prices

### v1

@@ 1 Introduction
Trustworthy electricity price forecasting (EPF) systems represent fundamental strategic tools for utilities, retailers, aggregators, and large consumers to achieve effective participation in liberalized energy markets. Beside being exploited to anticipate price movements and perform bidding strategies, EPF is a key enabler of further crucial decision-making stages [1], such as optimal generation and asset management [2], energy-aware planning and scheduling [3], etc. In particular, day-ahead EPF - i.e., the prediction of the hourly prices for the next day - is a challenging task subject of continuous interest in both academia and industry [4]. Compared to other traded commodities, electricity is still not economically storable on a large scale; thus, a constant balance between demand and supply is essential to achieve overall system stability. Hence, the price profiles in liberalized power markets often exhibit quite peculiar characteristics [5]. Complex relationships to conditioning variables (such as load demand, fuel costs, electricity production and weather conditions) are typically involved, including meshed short and long-term seasonalities, as well as sensible volatility, usually orders of magnitude larger than in other utility tradings [6], [7], [8], [9].
Critical deviations have been injected by the repercussions of the steep gas price variations on the power plants. The increasing penetration of renewable sources in the generation mix, fundamental to contrast the global warming, is introducing further short-term price fluctuations to be promptly tackled. Additional variables such as the CO2 certificate prices interact with the power price features [10]. More complex market dynamics to be tackled are those injected by the expanding competition between an increasing number of highly active participants pursuing novel business models [11],[12].
Overall, forecasting is becoming harder to perform than ever before, and advances in this field are strongly requested to address such an evolving context [4], [13].
@@ 1.1 Literature review
Over the years, a broad set of approaches have been investigated to perform EPF. A non-exhaustive list includes auto-regressive models and related extensions, exponential smoothing, generalized additive models, Gaussian Process, gradient boosting, neural networks, support vector machines, random forest, fuzzy logic, wavelet ensembles, and hybrid models. More detailed descriptions and comparisons of these techniques can be found in [14], [15],[1], and references therein. According to the aforementioned reviews, the EPF literature was primarily dedicated to simple statistical models and small neural network architectures until the early 2010s, mainly due to computational constraints of the technologies available. Besides, they proved effective in addressing the more stable price series occurring during that period.
@@
Following the demand for more enhanced modeling capabilities to properly tackle modern energy market dynamics, the EPF community has been investigating more flexible but computationally intensive machine learning approaches [16], [12], [17].
@@
Within this rapidly developing and fascinating field of research, much interest is being devoted to modern neural network architectures [18],[19]. Such research momentum is mainly propelled by the significant results achieved across a broad set of computer science applications, including natural language processing [20] and computer vision [21]. Several research studies have contributed to the exploration and development of neural networks (NN) in the EPF field (see e.g., [4], [22], [23] for recent reviews). To date, the most comprehensive experimental investigation of NN based approaches to EPF is provided by [24], including feed-forward, recurrent and convolutional models, as well as several more traditional statistical techniques (such as ARX, ARIMA-GARCH, etc.), primarily adopted in the literature. When tested on the EPEX (Belgium day-ahead market benchmark), a deep feed-forward architecture (labeled DNN hereafter) achieved the highest predictive accuracy, outperforming both the traditional methods and the best recurrent neural networks with LSTM and GRU cells.
@@
In continuation of the previous works, in [4],[25] authors have further compared the DNN architecture to the recently proposed LASSO Estimated AutoRegressive model, arguably the most accurate approach yet in the broad statistical family. By deploying an extensive automatic hyperparameter optimization and evaluating a set of benchmark EPF tasks covering five different day-ahead markets, they confirmed the performance gains provided by the more flexible DNN. Indeed, the latter paper has been entitled “the dawn of machine learning” for EPF. Following these results, several NN-based techniques have been explored in recent years with the aim of reducing prediction error (see e.g., [26] and references therein). In economic terms, even single percentage gains can lead to annual savings of up to millions of dollars for companies [13].
@@
Although most of the available studies address point predictions (e.g., mean or median values) of day-ahead hourly prices, an increasing research momentum targets NN based probabilistic EPF (PEPF). Specifically, PEPF is aimed to characterize the underlying uncertainty of the model in the target space, e.g., via prediction intervals, quantiles or full predictive distributions. This is essential for companies operating within power market of increasing volatility, e.g., to support effective risk optimization, stochastic optimization, what-if scenario analysis before trading, etc. Several recent studies investigate this topic from different perspectives, reporting both additional problems and opportunities insisting on the EPF task (see e.g., [27], [28], [29] and references therein for details). A through empirical comparison of deep learning architectures for multi-variate probabilistic energy forecasting is performed in [22], including DeepAR, DeepTCN, DSANet and LSTNet models.
@@
According to [30],[31], Quantile Regression Average (QRA) represents a reference benchmark to turn point NNs into a PEPF setup, considering the ranking in the GEFCom competition [1]. Distributional NNs [22] and Bayesian Deep learning [23] techniques have been proposed, approximating the conditional distribution of the hourly prices given the input features, through parametric forms. The Gaussian distribution is commonly employed for this purpose. However, it has been observed that the prices tend to assume complex patterns, including heterosckedasticity, sensible skewedeness and fat tails [30],[32]. Deep quantile regression (DQR) based methods have been introduced to support non-parametric uncertainty characterization (see e.g., [33] and references therein). Despite their asymptotic validity, they inherit potential quantile overfitting and overconfidence in finite samples, particularly on the extreme quantiles. Ensembles of distributional NNs (i.e., Deep Ensembles) parameterizing flexible Johnson’s SU distributions have been recently proposed for this purpose, outperforming the state of the art PEPF benchmarks including both conventional Gaussian forms and the widely applied quantile regression on NN ensembles [30]. Despite the relevant improvements, it has been observed that such models lack the required calibration capabilities - i.e., the statistical consistency between the probabilistic forecasts and the observations -, failing to pass the coverage tests at various hours on the prediction horizon.
@@
Conformal prediction (CP) provides a principled framework to attain distribution free finite sample marginal calibration guarantees. Pioneered in the early 2000s, CP has recently become subject of a renewed research momentum within the statistical and machine learning communities (see e.g., [34] and reference therein for a detailed review). However, CP has still attracted minor attention in the EPF field. To date, only two papers have investigated CP-based approaches for this purpose. A first empirical study has been performed by [31] where the conventional split CP is compared to a normalized absolute deviation on LASSO, KNN, SVM models, achieving promising results. Later, [35] implemented an adaptive conformal inference procedure on Random Forest based point predictors. As this presents a relatively novel and still understudied research direction, further extensions are expected from the community in the near future.
@@
By looking beyond the EPF literature, the last developments in the conformal prediction research domain provides a lot of compelling ideas to be further explored for achieving sharp and calibrated probabilistic forecasts (see e.g., [36] and references therein). To the best of our knowledge, the state-of-art neural network-based approaches for PEPF have not yet been compared with CP extensions to assess potential probabilistic performance gains.
@@ 1.2 Contributions and organization of the paper
Then, we propose a novel approach to PEPF, extending the Deep Ensembles based methods through conformal inference based techniques deployed within an on-line recalibration procedure. The developed method consists of the following major ingredients. First of all, we leverage the asymmetric CP formulation introduced by [37] for regression tasks, here developed through a daily recalibrated multi-horizon time series setup to support flexible step-wise compensations of each upper/lower band. Overall, this enhance local PIs reliability (i.e., sample-wise efficiency) beyond the CP marginal coverage. Beside, we deploy the adversarial conformal inference setup proposed in [38] for uncertainty quantification under non-exchangeable conditions, such as distribution shifts. Hence, the target prediction bands over the different coverage levels are dynamically adjusted by tracking the related quantiles of the conformity score sequences, incorporating the running sum of the miscoverage events to support a more stable long-run calibration. To estimate the quantiles to be calibrated, we explore both a DQR setup and samples from different parameterized distributional NNs to assess their potential impact on both sample efficiency and reliability. A uniform vincentization technique is exploited for ensembles aggregation, leading to sharper bounds than the alternative probability aggregation, beside marginalizing the different local minimizers reached by the training algorithm. Moreover, a post-hoc sorting operator has been included before combination to achieve conditional quantile non-crossing.
@@ 2 Methods
In this section, we deepen the developed PEPF approach. We start by providing a brief introduction to the general conformal inference framework. Then, we address the local adaptivity issue by deepening the conformalized quantile regression techniques. Subsequently, we describe the methods deployed to address the lack of robustness of CP under non-exchangeable conditions. Finally, we detail the prediction quantiles estimation procedures, the network archiectures employed, as well the deep ensemble combination techniques.
@@ 2.1 Conformal Prediction
$\mathbb{P}(y_{t}\in\underbrace{f(x_{t})\pm\mathcal{S}_{(\lceil(n+1)(1-\alpha)\rceil)}}_{C_{1-\alpha}(x_{t})})=\frac{\lceil(n+1)(1-\alpha)\rceil}{(n+1)}$ (1)
@@ 2.2 Improved local adaptivity by conformalized quantiles
As apparent from the Eq.1, the conventional CP based on absolute scores results in symmetric average bands around the NN predictions, thus lacking the capability to adapt the PIs width on simpler/harder test conditions (e.g., heteroscedastic). Since full conditional coverage (i.e., $\mathbb{P}(y_{n+1}\in C(x_{n+1})|x_{n+1}=\chi)\geq 1-\alpha$, for almost all $\chi$) is not reachable in finite samples under weak distribution-free assumption, the goal is to obtain sample-wise efficiency (local PIs adaptivity) with proven marginal validity.
@@
where $\mathcal{C}^{h}_{1-\alpha}(x_{t})$ represents (1-$\alpha$)-level PI at prediction step $h=[1,...,H]$ given the input features. $\small{{l}^{h}_{1-\alpha/2}\small}$ and ${u}^{h}_{1-\alpha/2}$ depict the (1-$\alpha/2$)-th empirical quantiles of ${q}^{h}_{\alpha/2}(x_{i})-y_{i}^{h}:i\in\mathcal{I}_{c_{t}}$ and $y_{i}^{h}-{q}^{h}_{1-\alpha/2}(x_{i}):i\in\mathcal{I}_{c_{t}}$ respectively, computed on the calibration subset $\mathcal{I}_{c_{t}}$ for test time $t$. $\{{q}^{h}_{{\gamma}}(x_{j})\}_{{\gamma}\in\Gamma},\>{q}^{h}_{{\gamma}}(x_{j})\leq{q}^{h}_{{\gamma}^{\prime}}(x_{j})\;\forall{\gamma}<{\gamma}^{\prime}$ represent the discrete set $\Gamma$ of the quantiles predicted by the NN ensemble (see Section 2.4), while $y_{t}^{{h}}$ are the true values. The PI bounds at specific coverage levels are straight derived as the predicted quantiles in the set $\Gamma$ are defined by balanced pairs (e.g., distribution deciles, percentiles, etc).
Thanks to such formulation, upper/lower quantiles can be specifically adjusted at each stage over the prediciton horizon, considering the related coverage on the calibration sets. Overall, this yields stronger coverage capabilities by enabling more efficient input features conditioned approximate PIs estimation beyond the CP marginal coverage. In fact, more flexibility in compensating the predicted bands is deemed useful for addressing the complex shapes typically involved in PEPF applications, such as sensible heteroskedasticity, skewness and fat tails [30]. Besides, the asymptotic validity supported by quantile regression NNs (i.e., DQR) is extended to reach finite samples coverage guarantees, while compensating for potential quantile overfitting and overconfidence.
@@ 2.3 Addressing non-exchangeable conditions
Three main classes of approaches have been recently investigated in the CP literature devoted to non-exchangeable time series tasks, namely sequential split CP (see e.g. [31]), bootstrapped estimators without underlying models refitting (as e.g., in [42],[43]) and adversarial CP settings (see e.g., [44],[35],[38]).
@@
The motivations behind such methodological choices are summarized hereafter. First of all, while in principle CP yields marginal coverage despite the accuracy of the underlying model, the latter impacts the local reliability of the prediction intervals [36]. Hence, updating the DE by exploiting the last observations can support probabilistic performances beside point accuracy. Besides, the ensemble combination (detailed in Section 2.4) provides a simple but efficient technique to get rid of poor local minimizers in the network parameter space, often leading to improved accuracy beyond single networks (see e.g., [4]). Despite more computationally demanding, online DE retraining is feasible for EPF applications (see e.g., [30]). For instance, it is commonly employed within the widely adopted Quantile Regression Averaging (QRA) technique, as well as for distributional NNs aggregations in the litarature. The investigation of conformalized quantile regression methods without DE recalibration (as e.g. in [43]) and further on-line adaptive CP wrappers (as e.g., [35]) is left to future works. Finally, the quantiles tracking and integration provides further mechanism to compensate coverage degradation and systematic errors occurring over the recalibration window, e.g., due to sensible short-term drifts. This support a principled long-run reliability, i.e., $\lim_{T\to+\infty}1/T\sum_{t=1}^{T}{1}\{y_{t}^{h}\not\in\mathcal{C}^{h}_{1-\alpha}(x_{t})\}=\alpha$, for each target coverage degree $1-\alpha$ (see [38] for further details on long-run coverage and related proofs). The exploration of the derivative action (i.e., the scorecasting component) is left to future works devoted to specific robust design for addressing the potential degrade in stability.

### v2

@@ 1 Introduction
Trustworthy electricity price forecasting (EPF) systems represent fundamental strategic tools for utilities, retailers, aggregators, and large consumers to achieve effective participation in liberalized energy markets. Beside being exploited to anticipate price movements and perform bidding strategies, EPF is a key enabler of further crucial decision-making stages [1], such as optimal generation and asset management [2] and energy-aware planning and scheduling [3]. In particular, hourly day-ahead EPF - i.e., the prediction of the hourly prices for the next day - is a challenging task subject of continuous interest in both academia and industry [4]. Compared to other traded commodities, electricity is still not economically storable on a large scale; thus, a constant balance between demand and supply is essential to achieve overall system stability. Hence, the price profiles in liberalized power markets often exhibit quite peculiar characteristics [5]. Complex relationships with conditioning variables (such as load demand, fuel costs, electricity production and weather conditions) are typically involved, encompassing meshed short and long-term seasonalities, as well as considerable volatility, usually orders of magnitude larger than in other utility trading contexts [6], [7], [8], [9].
Critical deviations have been injected by the repercussions of the steep gas price variations on the power plants. The increasing penetration of renewable sources in the generation mix, fundamental to contrast the global warming, is introducing further short-term price fluctuations to be promptly tackled. Additional variables such as the CO2 certificate prices interact with the power price features [10]. Additional complex market dynamics arise from the increasing competition among a growing number of highly active participants, each pursuing innovative business models [11],[12].
Overall, forecasting is becoming harder to perform than ever before, and advances in this field are strongly requested to cope with such an evolving context [4], [13].
@@ 1.1 Literature review
Over the years, a broad set of approaches have been investigated to perform EPF. A non-exhaustive list includes auto-regressive models and related extensions, exponential smoothing, generalized additive models, Gaussian Process, gradient boosting, neural networks, support vector machines, random forest, fuzzy logic, wavelet ensembles, and hybrid models. More detailed descriptions and comparisons of these techniques can be found in [14], [15],[1], and references therein. According to the aforementioned reviews, the EPF literature was primarily dedicated to simple statistical models and small neural network architectures until the early 2010s, mainly due to computational constraints of the technologies available. Besides, they proved effective in addressing the more stable price series occurring during that period [16].
@@
Following the demand for more enhanced modeling capabilities to properly tackle modern energy market dynamics, the EPF community has been investigated more flexible but computationally intensive machine learning approaches [17], [12], [18], [19].
@@
Within this rapidly developing and fascinating field of research, much interest is being devoted to modern neural network architectures [20],[21]. Such research momentum is mainly propelled by the significant results achieved across a broad set of computer science applications, including natural language processing [22] and computer vision [23]. Several research studies have contributed to the exploration and development of neural networks (NN) in the EPF field (see e.g., [4], [24], [25] for recent reviews). To date, the most comprehensive experimental investigation of NN based approaches to EPF is provided by [26], including feed-forward, recurrent and convolutional models, as well as several more traditional statistical techniques (such as ARX, ARIMA-GARCH, etc.), primarily adopted in the literature. When tested on the EPEX (Belgium day-ahead market benchmark), a deep feed-forward architecture (labeled DNN hereafter) achieved the highest predictive accuracy, outperforming both the traditional methods and the best recurrent neural networks with LSTM and GRU cells  [^1] LSTM: Long short-term memory, GRU: Gated recurrent unit.
@@
Building upon the previous works, in [4] and [27] the authors have conducted a comparison betwee the DNN architecture and the recently proposed LASSO Estimated AutoRegressive model, which is arguably the most accurate approach within the broader statistical family. By deploying an extensive automatic hyperparameter optimization and evaluating a set of benchmark EPF tasks covering five different day-ahead markets, they confirmed the performance gains provided by the more flexible DNN. Indeed, the latter paper has been entitled “the dawn of machine learning” for EPF. Following these results, several NN-based techniques have been explored in recent years with the aim of reducing prediction error (see e.g., [28] and references therein). In economic terms, even single percentage gains can lead to annual savings of up to millions of dollars for companies [13].
@@
Although most of the available studies address point predictions (e.g., mean or median values) of day-ahead hourly prices, an increasing research momentum targets NN-based probabilistic EPF (PEPF). Specifically, PEPF is aimed to characterize the underlying uncertainty of the model in the target space, e.g., via prediction intervals, quantiles or full predictive distributions. This is essential for companies operating within power markets showing increasing volatility, e.g., to support extremely relevant business tasks, such as effective risk optimization, stochastic optimization and what-if scenario analysis before trading. Several recent studies investigate this topic from different perspectives, reporting both additional problems and opportunities insisting on the EPF approach (see e.g., [29], [30], [31], [32] and references therein for details). An thorough empirical comparison of deep learning architectures for multivariate probabilistic energy forecasting is performed in [24], including DeepAR, DeepTCN, DSANet and LSTNet models.
According to several research works (see e.g., [1],[33], [34] and references therein), Quantile Regression Average (QRA) represents the reference benchmark to turn point NNs into a PEPF setup, following the ranking in the GEFCom competition  [^2] The Global Energy Forecasting Competition (GEFCom) began in August 2014, with a specific emphasis on probabilistic energy forecasting, including load, price, wind, and solar predictions. The price category drew interest from 287 participants globally, with the top-performing teams subsequently invited to contribute papers to a special 2016 edition of the International Journal of Forecasting [35].
@@
Distributional NNs [24] and Bayesian Deep learning [25] techniques have been proposed approximating the conditional distribution of the hourly prices given the input features, through parametric forms. The Gaussian distribution is commonly employed for this purpose. However, it has been observed that the prices tend to exhibit complex patterns, including heterosckedasticity, significant skewness and fat tails [33] and [36]. Deep quantile regression (DQR) based methods have been introduced to support non-parametric uncertainty characterization (see e.g., [37] and references therein). Despite their asymptotic properties, they inherit potential quantile overfitting and overconfidence on the extreme quantiles in finite samples. Hence, Ensembles of distributional NNs (i.e., Deep Ensembles) parameterizing flexible Johnson’s SU distributions have been recently proposed for this purpose, outperforming the state of the art PEPF benchmarks including both conventional Gaussian forms and the widely applied quantile regression on NN ensembles [33]. Despite the relevant improvements, it has been observed that such models lack the required calibration capabilities - i.e., the statistical consistency between the probabilistic forecasts and the observations - failing to pass the coverage tests at various hours on the prediction horizon.
@@
Conformal Prediction (CP) provides a principled framework to attain distribution-free finite sample marginal calibration guarantees. Pioneered in the early 2000s, CP has recently become subject of a renewed research interest within the statistical and machine learning communities (see e.g., [38] and reference therein for a detailed review). However, CP has so far attracted minor attention in the EPF field. To date, only two papers have investigated CP-based approaches for this purpose. A first empirical study has been performed by [34] where the conventional split CP is compared to a normalized absolute deviation on LASSO, KNN, SVM models, achieving promising results. Later, [39] implemented an adaptive conformal inference procedure on Random Forest based point predictors. As this presents a relatively novel and still understudied research direction, further extensions are expected from the community in the near future.
@@
By looking beyond the EPF literature, last developments in the Conformal Prediction research domain provides a lot of compelling ideas to be further explored for achieving sharp and calibrated probabilistic forecasts (see e.g., [40] and references therein). To the best of our knowledge, the state-of-the-art neural network-based approaches for PEPF have not yet been compared with CP extensions to assess potential probabilistic performance gains.
@@ 1.2 Contributions and organization of the paper
Then, we propose a novel approach to PEPF, extending the Deep Ensembles based methods through conformal inference based techniques deployed within an on-line recalibration procedure. The developed method consists of the following major ingredients. First of all, we leverage the asymmetric CP formulation introduced by [41] for regression tasks, here developed through a daily recalibrated multi-horizon time series setup to support flexible step-wise compensations of each upper/lower band. Overall, this enhance local PIs reliability (i.e., sample-wise efficiency) beyond the CP marginal coverage. Beside, we deploy the adversarial conformal inference setup proposed in [42] for uncertainty quantification under non-exchangeable conditions, such as distribution shifts. Hence, the target prediction bands over the different coverage levels are dynamically adjusted by tracking the related quantiles of the conformity score sequences, incorporating the running sum of the miscoverage events to support a stable long-run calibration. To estimate the quantiles to be calibrated, we explore both a DQR setup and samples from different parameterized distributional NNs to assess their potential impact on both sample efficiency and reliability. A uniform vincentization technique is exploited for ensembles aggregation, leading to sharper bounds than the alternative probability aggregation, beside marginalizing the different local minimizers reached by the training algorithm. Moreover, a post-hoc sorting operator is included before combination to achieve conditional quantile non-crossing.
We structured an open GitHub repository with the datasets and code to reproduce the experiments, as well as to support the integration and comparison of further datasets and PEPF techniques  [^3] https://github.com/bruale/PefCodeBench.
@@ 2 Methods
In this section, we deepen the developed PEPF approach. We start by providing a brief introduction to the general conformal inference framework. Then, we address the local adaptivity issue by deepening the conformalized quantile regression technique. Subsequently, we describe the methods deployed to address the lack of robustness of CP under non-exchangeable conditions. Finally, we detail the prediction quantiles estimation procedures, the network archiectures employed, as well the deep ensemble combination techniques.
@@ 2.1 Conformal Prediction
$\mathbb{P}(y_{t}\in\underbrace{f(x_{t})\pm\mathcal{S}_{(\lceil(n+1)(1-\alpha)\rceil)}}_{C_{1-\alpha}(x_{t})})=\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ (1)
@@ 2.2 Improved local adaptivity by conformalized quantiles
As apparent from the Eq.1, the conventional CP based on absolute scores results in symmetric average bands around the NN predictions, thus lacking the capability to adapt the PIs width on simpler/harder test conditions (e.g., heteroscedasticity). Since full conditional coverage (i.e., $\mathbb{P}(y_{t}\in C(x_{t})|x_{t}=\chi)\geq 1-\alpha$, for almost all $\chi$) is not reachable in finite samples under weak distribution-free assumption, the goal is to obtain sample-wise efficiency (local PIs adaptivity) with proven marginal validity.
@@
where $\mathcal{C}^{h}_{1-\alpha}(x_{t})$ represents (1-$\alpha$)-level PI at prediction step $h=[1,...,H]$ given the input features. $\small{{l}^{h}_{1-\alpha/2}\small}$ and ${u}^{h}_{1-\alpha/2}$ depict the (1-$\alpha/2$)-th empirical quantiles of ${q}^{h}_{\alpha/2}(x_{i})-y_{i}^{h}:i\in\mathcal{I}_{c_{t}}$ and $y_{i}^{h}-{q}^{h}_{1-\alpha/2}(x_{i}):i\in\mathcal{I}_{c_{t}}$ respectively, computed on the calibration subset $\mathcal{I}_{c_{t}}$ for test time $t$. $\{{q}^{h}_{{\gamma}}(x_{j})\}_{{\gamma}\in\Gamma},\>{q}^{h}_{{\gamma}}(x_{j})\leq{q}^{h}_{{\gamma}^{\prime}}(x_{j})\;\forall{\gamma}<{\gamma}^{\prime}$ represent the discrete set $\Gamma$ of the quantiles predicted by the NN ensemble (see section 2.4), while $y_{t}^{{h}}$ are the observed values. The PI bounds at specific coverage levels are straight derived as the predicted quantiles in the set $\Gamma$ are defined by balanced pairs (e.g., distribution deciles, percentiles, etc).
Thanks to such formulation, upper/lower quantiles can be specifically adjusted at each stage over the prediction horizon, considering the related coverage on the calibration sets. Overall, this yields stronger coverage capabilities by enabling more efficient input features conditioned approximate PIs estimation beyond the CP marginal coverage. In fact, more flexibility in compensating for the predicted bands is deemed useful for addressing the complex shapes typically involved in PEPF applications, such as heteroscedasticity, skewness and fat tails [33]. Besides, the asymptotic validity supported by quantile regression NNs (i.e., DQR) is extended to reach finite samples coverage guarantees, while compensating for potential quantile overfitting and overconfidence.
@@ 2.3 Addressing non-exchangeable conditions
Three main classes of approaches have been recently investigated in the CP literature devoted to non-exchangeable time series tasks, namely online sequential split CP (see e.g. [34]), bootstrapped estimators without underlying models refitting (as e.g., in [46],[47]) and adversarial CP settings (see e.g., [48],[39],[42]).
@@
The motivations behind such methodological choices are summarized hereafter. First of all, while in principle CP yields marginal coverage despite the accuracy of the underlying model, the latter impacts the local reliability of the prediction intervals [40]. Hence, updating the DE by exploiting the last observations can support probabilistic performances beside point accuracy. Besides, the ensemble combination (detailed in section 2.4) provides a simple but efficient technique to get rid of poor local minimizers in the network parameter space (see section), often leading to improved accuracy beyond single networks (see e.g., [4]). Despite more computationally demanding, online DE retraining is feasible for EPF applications (see e.g., [33]). For instance, it is commonly employed within the widely adopted Quantile Regression Averaging (QRA) technique, as well as for distributional NNs aggregations in the litarature. The investigation of conformalized quantile regression methods without DE recalibration (as e.g. in [47]) and further on-line adaptive CP wrappers (as e.g., [39]) is left to future works. Finally, the quantiles tracking and integration provides further mechanism to compensate coverage degradation and systematic errors occurring over the recalibration window, e.g., due to sensible short-term drifts. This support a principled long-run reliability, i.e., $\lim_{T\to+\infty}1/T\sum_{t=1}^{T}1\{y_{t}^{h}\not\in\mathcal{C}^{h}_{1-\alpha}(x_{t})\}=\alpha$, for each target coverage degree $1-\alpha$ (see [42] for further details on long-run coverage and related proofs). The exploration of the derivative action (i.e., the scorecasting component) is left to future works devoted to specific robust design for addressing the potential degrade in stability.